OpenStudy (anonymous):

Emma and Rita are at a florist shop. They are planning to make bouquets of flowers for their grandmothers out of roses and carnations. The cost for a single rose is r, and the cost for a single carnation is c. Emma will use 6 carnations and 4 roses for her bouquet, and Rita will use 5 carnations and 6 roses. Write expressions that represent the total cost of Emma’s bouquet, the total cost of Rita’s bouquet, and the total cost for both bouquets combined.

OpenStudy (anonymous):

@david111

OpenStudy (anonymous):

11x + 10r = x

OpenStudy (anonymous):

6c + 4r = Emma's bouquet 5c + 6r = Rita's bouquet Then .. 11c + 10r = both bouquets

OpenStudy (anonymous):

thank you for your assistance.

OpenStudy (anonymous):

No problm

OpenStudy (anonymous):

*No problem